# 11footnotetext: Corresponding author

### v1

@@ Abstract
Collaborative perception improves task performance by expanding the perception range through information sharing among agents. . Immutable heterogeneity poses a significant challenge in collaborative perception, as participating agents may employ different and fixed perception models. This leads to domain gaps in the intermediate features shared among agents, consequently degrading collaborative performance. Aligning the features of all agents to a common representation can eliminate domain gaps with low training cost. However, in existing methods, the common representation is designated as the representation of a specific agent, making it difficult for agents with significant domain discrepancies from this specific agent to achieve proper alignment. This paper proposes NegoCollab, a heterogeneous collaboration method based on the negotiated common representation. It introduces a negotiator during training to derive the common representation from the local representations of each modality’s agent, effectively reducing the inherent domain gap with the various local representations. In NegoCollab, the mutual transformation of features between the local representation space and the common representation space is achieved by a pair of sender and receiver. To better align local representations to the common representation containing multimodal information, we introduce structural alignment loss and pragmatic alignment loss in addition to the distribution alignment loss to supervise the training. This enables the knowledge in the common representation to be fully distilled into the sender. The experimental results demonstrate that NegoCollab significantly outperforms existing methods in common representation-based collaboration approaches. The mechanism of obtaining common representations through negotiation provides a more reliable and flexible option for common representations in heterogeneous collaborative perception.
@@ 1 Introduction
Collaborative perception has gained significant attention in recent years. By sharing intermediate features among agents, it expands the perception range and provides more supporting information for downstream tasks. In autonomous driving, collaborative perception enables vehicles to detect obstacles in blind spots, thereby preventing traffic accidents and effectively enhancing driving safety. Heterogeneity is one of the key challenges in practical applications of collaborative perception xu2023mdpa; lu2024heal; gao2025stamp. When there are differences in sensors and perception models among collaborating agents, it creates domain gaps in the shared intermediate features. This prevents effective fusion of features from heterogeneous agents and consequently degrades collaborative performance.
Current research on heterogeneity issues includes approaches that achieve heterogeneous collaboration by retraining specialized collaborative modules xiang2023hmvit or sharing partial networks in model lu2024heal. However, in practical deployment, perception model are crucial for autonomous driving safety and tightly coupled with downstream tasks, making it difficult to replace or retrain. These limitations lead to the challenge of immutable heterogeneous collaborative perception xia2024polyinter. To address this issue, methods like xu2023mdpa; pnpda; xia2024polyinter employ domain adapters or polymorphic prompts to eliminate domain gaps through one-to-one adaptation for heterogeneous agents, as is shown in Figure 1a, requiring only single-step feature transformation but incurring higher training costs. Alternatively, gao2025stamp aligns the representations of each modality’s agent to a common representation by training a pair of adapter and reverter, which has low training cost. However, since the common representation is designated as the representation of a specific agent, as is shown in Figure 1b, alignment becomes difficult to achieve when there exists a large domain gap among the representations of other agents and the designated agent.
@@ 2.1 Collaborative Perception
In recent years, collaborative perception has attracted widespread attention due to its potential to enhance autonomous driving safety. By sharing perception data among agents—including raw sensor data rauch2012car2x; luo2023edgecooper; liu2024v2x-pc, intermediate features wang2020v2vnet; li2021disconet-learning; fcooper; hu2022where2comm, and detection results xu2023Model-agnostic; Rawashdeh2018CoAD, collaborative perception effectively expands the perception range and overcomes blind spots and occlusion issues inherent in single-agent perception. However, in real-world scenarios, collaborative perception faces multiple challenges including: limited communication bandwidth hu2022where2comm; hu2023coca; hu2024codefilling, location noise lu2023coalign; lei2024freealign, communication delay and computation asynchronously lei2022syncnet; wei2024cobevflow, communication interruptions ren2024incop, heterogeneity xu2023mdpa; xiang2023hmvit; lu2024heal; pnpda; gao2025stamp; xia2024polyinter, security and privacy concerns li2023robosac; zhao2023malicious-det, and simulation-to-real generalization issues kong2023dusa; wei2024chatsim-auto, all of which pose challenges to collaboration. This paper focuses on the heterogeneity challenge in collaborative perception, proposing a negotiated common representation-based approach to achieve common representation-based heterogeneous collaboration.
@@ 2.2 Multi-modal Representation Learning
Multi-modal representation learning manzoor2023mmsurvey enables information fusion and transformation across different modalities (e.g., images, LiDAR point clouds, text, speech) by learning a shared representation space. In autonomous driving, approaches like zhang2025sparselif; liu2023bevfusion; lu2024heal employ network designs such as sparse transformers and feature pyramids to learn fused multi-modal representations from LiDAR point clouds and camera images, significantly enhancing vehicles’ environmental perception capabilities. Knowledge distillation serves as a common method for cross-modal knowledge transfer, approaches like zhou2023unidistill; wang2024distilvpr; chen2022bevdistill apply various distillation losses, including dense distillation loss, relative relation distillation loss, and response distillation loss, between multi-modal features to achieve mutual enhancement of multi-modal information, thereby improving task performance. This paper generates the common representation from the local representations of each modality using a feature pyramid network, while introduces a multi-dimensional alignment loss composed of distribution alignment loss, structural alignment loss, and pragmatic alignment loss during training to enable more effective alignment of local representations to the multi-modal common representation.
@@ 3.1.1 Sender
The sender’s role is to transform features from the local representation space to the common representation space, consisting of two modules: recombiner and aligner. The recombiner employs a ConvNeXtLiu_2022_convnext structure to enhance local features beneficial for collaboration. It also includes a size-channel alignment module to adjust the dimensions and channels of local features to standard settings. The aligner uses a fused axial attention xu2022cobevt to capture both global and local dependencies within features, thereby mapping features from the local representation space to the common representation space.
@@ 3.2.3 Multi-dimensional Information Alignment
where $L_{focal}$ is the focal loss lin2017focal.
@@ 4.1 Settings
Align to
Common
NegoCollab 0.872 0.911 0.512 0.745 0.765 0.854 0.319 0.555
@@
Align to
Common
MPDA-P 0.467 0.334 0.258 0.211
@@ 4.2 Quantitative Analysis
Performance of heterogeneous collaboration. We evaluated each method on the OPV2V-H lu2024heal, V2V4Real xu2023v2v4real, and DAIR-V2X yu2022dair datasets, as shown in Table 1 and Table 2. Since the common representation of MPDA-P, PnPDA-P, and STAMP are all derived from the single-modality protocol agent, for fair comparison, we implement NegoCollab-P, which derives the common representation from the protocol agent. In Table 1, the columns m1m2, m1m3, m2m4, and m1m2m3m4 correspond to the performance of: initial alliance agents, heterogeneous LiDAR agents, heterogeneous camera agents, and all agent types collaborative, respectively. The results demonstrate that among heterogeneous collaboration methods based on common representation, NegoCollab achieves the best performance in all test conditions. Compared with one-to-one adaptation methods, NegoCollab also maintains optimal collaborative performance when agents m1 and m2 within the initial alliance collaborated. For collaboration with new agents m3 and m4, although m3 and m4 did not participate in the negotiation process of the common representation, their collaborative performance is slightly lower than that of one-to-one adaptation methods, but still achieves competitive results. This strongly demonstrates NegoCollab’s superior performance and the excellent adaptability of the common representation to new agents. Additionally, the results in Table 2 show that NegoCollab also has excellent heterogeneous collaboration performance in real-world environments.
@@
Comparison of domain gaps. To validate the effectiveness of the negotiator in reducing domain gaps, we employ KL divergence kullback1951kldivergence to measure the domain gap between common representation and local representations of each modality across different methods. Comparision are illustrated in Figure 4. Since MPDA-P, PnPDA-P, and STAMP all use the representation of the protocol agent as the common representation, they are aggregated as the ’Baseline’ in the figure. It can be seen that the domain gap between the common representation generated by the negotiator and each local representation is significantly reduced. Compared to the method of directly designating the representation of the protocol agent as the common representation, the domain gap measured by KL divergence is reduced by an average of approximately 93.5
@@ 4.3 Ablation Study
Negotiating from different initial alliances. In practical applications, heterogeneous agents form multiple collaborative groups based on collaboration needs gao2025stamp, using different common representations for information sharing within each group. NegoCollab’s negotiation-based mechanism enables the free selection of agents from a collaborative group to negotiate the common representation, thereby providing more diverse and reliable common representation. To further explore how to negotiate a better common representation, we investigate the impact of common representations negotiated from different initial alliances on collaborative performance. Two key observations are summarized. with detailed content and experimental results provided in the Appendix.
@@ A.1 Dataset
OPV2V-H. OPV2V-H lu2024heal dataset contains 73 scenes covering 6 road types across 9 cities. Each Connected Autonomous Vehicle(CAV) in the scenes is equipped with one 16-channel, one 32-channel, and one 64-channel LiDAR, along with 4 monocular cameras and 4 depth cameras. The dataset comprises 36K frames of LiDAR point clouds, 12K frames of RGB camera images, 12K frames of depth camera images, and 230K annotated 3D bounding boxes.
DAIR-V2X. DAIR-V2X yu2022dair is a real-world collaborative perception dataset. The dataset has 9K frames featuring one vehicle and one roadside unit (RSU), both equipped with a LiDAR and a 1920x1080 camera. RSU’ LiDAR is 300-channel while the vehicle’s is 40-channel.
V2V4Real. V2V4Real xu2023v2v4real is a real-world Vehicle-to-Vehicle (V2V) cooperative perception dataset. The dataset includes 20,000 LiDAR scans and 240,000 annotated 3D bounding boxes across five vehicle classes. It supports benchmarks for three key task: 3D object detection, object tracking, and Sim2Real domain adaptation-enabling evaluation with state-of-the-art models.
@@ B.3 Component Ablation
Table 8: Component ablation study. The collaborating agents are m1 and m2, and the results are the performance without downstream collaborative task adaptation. The component name in bold in the settings column indicates the default configuration. Column corresponding to #Params# shows the number of parameters when the module uses the corresponding configuration. ‘M’ standing for ‘MB’. “ResMlp” is a network with a multi-layer perceptron as its backbone. FANetyoung2022fanet featuring an encoder-decoder structure, which can be used to adjust the feature space.

### v2

@@ Abstract
Collaborative perception improves task performance by expanding the perception range through information sharing among agents. Immutable heterogeneity poses a significant challenge in collaborative perception, as participating agents may employ different and fixed perception models. This leads to domain gaps in the intermediate features shared among agents, consequently degrading collaborative performance. Aligning the features of all agents to a common representation can eliminate domain gaps with low training cost. However, in existing methods, the common representation is designated as the representation of a specific agent, making it difficult for agents with significant domain discrepancies from this specific agent to achieve proper alignment. This paper proposes NegoCollab, a heterogeneous collaboration method based on the negotiated common representation. It introduces a negotiator during training to derive the common representation from the local representations of each modality’s agent, effectively reducing the inherent domain gap with the various local representations. In NegoCollab, the mutual transformation of features between the local representation space and the common representation space is achieved by a pair of sender and receiver. To better align local representations to the common representation containing multimodal information, we introduce structural alignment loss and pragmatic alignment loss in addition to the distribution alignment loss to supervise the training. This enables the knowledge in the common representation to be fully distilled into the sender. The experimental results demonstrate that NegoCollab significantly outperforms existing methods in common representation-based collaboration approaches. The mechanism of obtaining common representations through negotiation provides a more reliable and flexible option for common representations in heterogeneous collaborative perception.
@@ 1 Introduction
Collaborative perception has gained significant attention in recent years. By sharing intermediate features among agents, it expands the perception range and provides more supporting information for downstream tasks. In autonomous driving, collaborative perception enables vehicles to detect obstacles in blind spots, thereby preventing traffic accidents and effectively enhancing driving safety. Heterogeneity is one of the key challenges in practical applications of collaborative perception Xu et al. (2023b); Lu et al. (2024); Gao et al. (2025). When there are differences in sensors and perception models among collaborating agents, it creates domain gaps in the shared intermediate features. This prevents effective fusion of features from heterogeneous agents and consequently degrades collaborative performance.
Current research on heterogeneity issues includes approaches that achieve heterogeneous collaboration by retraining specialized collaborative modules Xiang et al. (2023) or sharing partial networks in model Lu et al. (2024). However, in practical deployment, perception model are crucial for autonomous driving safety and tightly coupled with downstream tasks, making it difficult to replace or retrain. These limitations lead to the challenge of immutable heterogeneous collaborative perception Xia et al. (2024). To address this issue, methods like Xu et al. (2023b); Luo et al. ; Xia et al. (2024) employ domain adapters or polymorphic prompts to eliminate domain gaps through one-to-one adaptation for heterogeneous agents, as is shown in Figure 1a, requiring only single-step feature transformation but incurring higher training costs. Alternatively, Gao et al. (2025) aligns the representations of each modality’s agent to a common representation by training a pair of adapter and reverter, which has low training cost. However, since the common representation is designated as the representation of a specific agent, as is shown in Figure 1b, alignment becomes difficult to achieve when there exists a large domain gap among the representations of other agents and the designated agent.
@@ 2.1 Collaborative Perception
In recent years, collaborative perception has attracted widespread attention due to its potential to enhance autonomous driving safety. By sharing perception data among agents—including raw sensor data Rauch et al. (2012); Luo et al. (2023); Liu et al. (2024), intermediate features Wang et al. (2020); Li et al. (2021); Chen et al. (2019); Hu et al. (2022), and detection results Xu et al. (2023a); Rawashdeh and Wang (2018), collaborative perception effectively expands the perception range and overcomes blind spots and occlusion issues inherent in single-agent perception. However, in real-world scenarios, collaborative perception faces multiple challenges including: limited communication bandwidth Hu et al. (2022, 2023, 2024), location noise Lu et al. (2023); Lei et al. (2024), communication delay and computation asynchronously Lei et al. (2022); Wei et al. (2024a), communication interruptions Ren et al. (2024), heterogeneity Xu et al. (2023b); Xiang et al. (2023); Lu et al. (2024); Luo et al. ; Gao et al. (2025); Xia et al. (2024), security and privacy concerns Li et al. (2023); Zhao et al. (2023), and simulation-to-real generalization issues Kong et al. (2023); Wei et al. (2024b), all of which pose challenges to collaboration. This paper focuses on the heterogeneity challenge in collaborative perception, proposing a negotiated common representation-based approach to achieve common representation-based heterogeneous collaboration.
@@ 2.2 Multi-modal Representation Learning
Multi-modal representation learning Manzoor et al. (2023) enables information fusion and transformation across different modalities (e.g., images, LiDAR point clouds, text, speech) by learning a shared representation space. In autonomous driving, approaches like Zhang et al. (2025); Liu et al. (2023); Lu et al. (2024) employ network designs such as sparse transformers and feature pyramids to learn fused multi-modal representations from LiDAR point clouds and camera images, significantly enhancing vehicles’ environmental perception capabilities. Knowledge distillation serves as a common method for cross-modal knowledge transfer, approaches like Zhou et al. (2023); Wang et al. (2024); Chen et al. (2022) apply various distillation losses, including dense distillation loss, relative relation distillation loss, and response distillation loss, between multi-modal features to achieve mutual enhancement of multi-modal information, thereby improving task performance. This paper generates the common representation from the local representations of each modality using a feature pyramid network, while introduces a multi-dimensional alignment loss composed of distribution alignment loss, structural alignment loss, and pragmatic alignment loss during training to enable more effective alignment of local representations to the multi-modal common representation.
@@ 3.1.1 Sender
The sender’s role is to transform features from the local representation space to the common representation space, consisting of two modules: recombiner and aligner. The recombiner employs a ConvNeXtLiu et al. (2022) structure to enhance local features beneficial for collaboration. It also includes a size-channel alignment module to adjust the dimensions and channels of local features to standard settings. The aligner uses a fused axial attention Xu et al. (2022) to capture both global and local dependencies within features, thereby mapping features from the local representation space to the common representation space.
@@ 3.2.3 Multi-dimensional Information Alignment
where $L_{focal}$ is the focal loss Lin et al. (2017).
@@ 4.1 Settings
Align to Common NegoCollab 0.872 0.911 0.512 0.745 0.765 0.854 0.319 0.555
@@
Align to Common MPDA-P 0.467 0.334 0.258 0.211
@@ 4.2 Quantitative Analysis
Performance of heterogeneous collaboration. We evaluated each method on the OPV2V-H Lu et al. (2024), V2V4Real Xu et al. (2023c), and DAIR-V2X Yu et al. (2022) datasets, as shown in Table 1 and Table 2. Since the common representation of MPDA-P, PnPDA-P, and STAMP are all derived from the single-modality protocol agent, for fair comparison, we implement NegoCollab-P, which derives the common representation from the protocol agent. In Table 1, the columns m1m2, m1m3, m2m4, and m1m2m3m4 correspond to the performance of: initial alliance agents, heterogeneous LiDAR agents, heterogeneous camera agents, and all agent types collaborative, respectively. The results demonstrate that among heterogeneous collaboration methods based on common representation, NegoCollab achieves the best performance in all test conditions. Compared with one-to-one adaptation methods, NegoCollab also maintains optimal collaborative performance when agents m1 and m2 within the initial alliance collaborated. For collaboration with new agents m3 and m4, although m3 and m4 did not participate in the negotiation process of the common representation, their collaborative performance is slightly lower than that of one-to-one adaptation methods, but still achieves competitive results. This strongly demonstrates NegoCollab’s superior performance and the excellent adaptability of the common representation to new agents. Additionally, the results in Table 2 show that NegoCollab also has excellent heterogeneous collaboration performance in real-world environments.
@@
Comparison of domain gaps. To validate the effectiveness of the negotiator in reducing domain gaps, we employ KL divergence Kullback and Leibler (1951) to measure the domain gap between common representation and local representations of each modality across different methods. Comparision are illustrated in Figure 4. Since MPDA-P, PnPDA-P, and STAMP all use the representation of the protocol agent as the common representation, they are aggregated as the ’Baseline’ in the figure. It can be seen that the domain gap between the common representation generated by the negotiator and each local representation is significantly reduced. Compared to the method of directly designating the representation of the protocol agent as the common representation, the domain gap measured by KL divergence is reduced by an average of approximately 93.5
@@ 4.3 Ablation Study
Negotiating from different initial alliances. In practical applications, heterogeneous agents form multiple collaborative groups based on collaboration needs Gao et al. (2025), using different common representations for information sharing within each group. NegoCollab’s negotiation-based mechanism enables the free selection of agents from a collaborative group to negotiate the common representation, thereby providing more diverse and reliable common representation. To further explore how to negotiate a better common representation, we investigate the impact of common representations negotiated from different initial alliances on collaborative performance. Two key observations are summarized. with detailed content and experimental results provided in the Appendix.
@@ A.1 Dataset
OPV2V-H. OPV2V-H Lu et al. (2024) dataset contains 73 scenes covering 6 road types across 9 cities. Each Connected Autonomous Vehicle(CAV) in the scenes is equipped with one 16-channel, one 32-channel, and one 64-channel LiDAR, along with 4 monocular cameras and 4 depth cameras. The dataset comprises 36K frames of LiDAR point clouds, 12K frames of RGB camera images, 12K frames of depth camera images, and 230K annotated 3D bounding boxes.
DAIR-V2X. DAIR-V2X Yu et al. (2022) is a real-world collaborative perception dataset. The dataset has 9K frames featuring one vehicle and one roadside unit (RSU), both equipped with a LiDAR and a 1920x1080 camera. RSU’ LiDAR is 300-channel while the vehicle’s is 40-channel.
V2V4Real. V2V4Real Xu et al. (2023c) is a real-world Vehicle-to-Vehicle (V2V) cooperative perception dataset. The dataset includes 20,000 LiDAR scans and 240,000 annotated 3D bounding boxes across five vehicle classes. It supports benchmarks for three key task: 3D object detection, object tracking, and Sim2Real domain adaptation-enabling evaluation with state-of-the-art models.
@@ B.3 Component Ablation
Table 8: Component ablation study. The collaborating agents are m1 and m2, and the results are the performance without downstream collaborative task adaptation. The component name in bold in the settings column indicates the default configuration. Column corresponding to #Params# shows the number of parameters when the module uses the corresponding configuration. ‘M’ standing for ‘MB’. “ResMlp” is a network with a multi-layer perceptron as its backbone. FANetYoung et al. (2022) featuring an encoder-decoder structure, which can be used to adjust the feature space.